# Characteristic initial value problems for the Einstein-Maxwell-scalar field equations in spherical symmetry

### v2

@@ III.2 Local initial-boundary formulation at a world line
In spherical symmetry there exists a designated worldline, which is the geodesic tracing the centers of spherical symmetry Synge (1960). We denote this geodesic with $c(w)$ and choose $w$ as an affine parameter for its parameterisation. The tangent vector $c^{a}\partial_{a}$ is a unit timelike vector and $w$ measures the proper time along $c(w)$. This proper time is assumed to be the proper time of a freely falling Fermi observer along $c(w)$ Manasse and Misner (1963); Poisson (2009). Following Madler and Muller (2013), in a convex normal neighborhood of $c(w)$ we extend the definition of $w$ be constant along null cones emanating from $c(w)$. The surface forming rays of the $w=const$ null cones are labeled by the angular coordinates $x^{A}$ within each null cone. The null condition on $w=const$ implies $g^{ww}=0$, since $g^{ab}\nabla_{a}w\nabla_{b}w=0$. The co-vector $\ell_{a}=\epsilon\nabla_{a}w$ defines by metric duality the tangent vectors $\ell^{a}=\epsilon g^{ab}\nabla_{b}w=\epsilon g^{aw}$ of the surface forming null rays. The null rays are labeled by the angular coordinates, $\ell^{a}\nabla_{a}x^{A}=0$ and therefore $g^{wA}=0$.
By these coordinate conditions, only the coordinate $x^{1}=\lambda$ varies along the rays of a given null cone $w=const$. Since the cones emanate from $c(w)$, we take the origin $\lambda=0$ to be on $c(w)$. We normalize the parameter $\lambda$ along the rays such that it is an affine parameter with the normalization determined at $c(w)$. In particular, $c^{a}\nabla_{a}\lambda=\ell^{a}\nabla_{a}\lambda=\epsilon g^{w\lambda}=1$ with $|\epsilon|=1$ at $c(w)$ so that $\epsilon=-1$ designates that the respective null cone $w=const$ are outgoing null cones with null rays pointing into the future of $c(w)$ and $\epsilon=1$ are ingoing null cones pointing into the past of $c(w)$.
By these conditions, the metric near $c(w)$ has the form [^4] RefMadler and Muller (2013) uses a reversed nomenclature for the binary parameter $\epsilon$. Madler and Muller (2013)
@@
The supplementary equation (II) holds at $c(w)$ due to the regularity conditions. Eq. (63) and (18) imply
$Q_{,w}=O(\lambda^{2})\;\;\Rightarrow\;\;Q(w,0)=Q_{,\lambda}(w,0)=0$ (65)
@@
where the subscript $(k)$ in parenthesis indicate that the symbol carrying the subindex is the $\lambda^{k}$ coefficient function depending on $w$ of the Taylor series of the respective field evaluated at $\lambda=0$.
The scalar field expansion (66) and (29h) give
@@
Insertion of these data into (29) and integration gives
@@
We can see that the time evolution of the coefficient $\Phi_{(i)}$ along $c(w)$ is driven by a function depending on the expansion coefficients containing all expansion coefficients up to $\Phi_{(i+1)}$. This is behavior is a result of the choice of $w$ being constant along a null hypersurface. The hierarchical structure of (69) is the scalar field equivalent to the similar structure in the gravitational case Madler and Muller (2013); Mädler (2013) for gravitational waves. Since $\Phi_{(1)}=\lim_{\lambda\rightarrow 0}\Phi_{,\lambda}$ and $\Phi_{(2)}=\lim_{\lambda\rightarrow 0}\Phi_{,\lambda\lambda}$, we learn from (68) that the evolution of the initial data $\Phi$ along $c(w)$ and in its neighbourhood is completely determined by the data on a cone $w=const$, provided the data and its derivatives are finite in the limit $\lambda\rightarrow 0$. Indeed it is not difficult to derive similar equations like in (69) for $\Phi_{(k)}$ with $k\geq 2$ while using an appropriate higher order expansion than (66) and integrating (29).
From such an exercise, we can deduce the following local characteristic initial-boundary value problem at the vertices of null cones emanating from a geodesic world line: Suppose, the complex scalar field $\Phi(w,\lambda)$ on a null cone $w=0$ has a regular Taylor series expansion,
@@
evaluated at the world line, the boundary conditions of the fields are (63) and (64), $\alpha_{0}$ is an arbitrary function of $w$ for $w>0$, and the gauge charge $q$ is provided, then the metric, the electromagnetic potential $\alpha$ and the charge function $Q$ can be completely calculated on the cone $w=0$ using the hypersurface equations (29a)-(29f). Their local solution at $\lambda=0$ is of the form (68a), (68b), (68c) or (68f). The evolution equations (29g) determine the time derivative of the Taylor coefficients of the initial data, as seen in e.g. (69).
We remark, that this regularity assumption at $w=0$ does not imply that the data necessarily remain regular for $w>0$. If $r\rightarrow 0$ for $\lambda\rightarrow\lambda_{C}>0$ a (massive) caustic forms at finite value for the affine parameter. This caustic formation for $w>0$ is in general driven by the value of norm squared, $|\Phi_{,\lambda}|^{2}$, of the derivative of the scalar field. An example for caustic initical data for a real scalar field was recently given in Mädler et al. (2024), where the data were employed in an evolution through an event horizon of a supercritical scalar field. The thus formulated initial value problem in spherical symmetry and with a scalar field is an analogue to the gravitational one of Friedrich Friedrich (1986).
@@ III.3 Initial-boundary value problem at a null hypersurface
Consider a family of null hypersurfaces $w=const$ designated by $\mathscr{N}^{\epsilon}(w)$, hereafter. The generators $\ell^{a}\partial_{a}$ of these null hypersurface are affinely parameterized with the affine parameter $\lambda$. We choose the origin, $\lambda=0$, of the affine parameter to be at a null hypersurface $\mathcal{B}^{\epsilon}$. The surface $\mathcal{B}^{\epsilon}$ is a null boundary for every null hypersurface $w=const$ of $\mathscr{N}^{\epsilon}$ and is the boundary surface for the double null CIBVPs to be presented, see Fig.(1).
Figure 1: A family of null hypersurfaces $\mathscr{N}^{\epsilon}(w)$ with corresponding boundary null hypersurface $\mathcal{B}^{\epsilon}(w)$ is shown for $\epsilon=-1$ (left) and $\epsilon=-1$ (right). The former is a characteristic initial value problem with the initial data surface $\mathscr{N}^{-}(0)$ and the latter gives rise to a characteristic past value problem with the past data on $\mathscr{N}^{+}(0)$. The common intersection between $\mathscr{N}^{\epsilon}(0)$ and $\mathcal{B}^{\epsilon}(w)$ is $\Sigma(0)$.
@@
This implies $V(w,0)=0$ for all values of $w$ and where $\lambda=0$. Consequently, $V=O(\lambda)$ near $\mathcal{B}^{\epsilon}$ on $\mathscr{N}^{\epsilon}$. For the conformal factor $r$ there are two possibilities regarding its evaluation on $\mathcal{B}^{\epsilon}$: i) $r(w,0)=0$ and ii) $r(w,0)\neq 0$.
Physically these cases mean that the null hypersurface $\mathcal{B}^{\epsilon}$ has either a trivial (case (i)) or non-trivial Misner-Sharp mass (case (ii)), which becomes clear considering (21) in the limit $\lambda\rightarrow\infty$
@@ III.3.1 Initial boundary formulation at a null hypersurface with $r(w,0)=0$
If $V(w,0)=r(w,0)=0$, the line element near the origin of the affine parameter $\lambda$ can be written like
@@
Consequently, we can set wlog $r_{(1)}=1$ which is assumed herefter in this subsection. With the assumption of a regular expansion of the scalar field like in (66), we find that $Z(w,0)=\mathcal{L}(w,0)=0$. The residual gauge freedom of the electromagnetic field $\alpha$, allows us to set $\alpha(w,0)=0$. In addition, the supplementary equation (18) implies $Q(w,0)=const$ for $\lambda=0$ and we may choose $Q(w,0)=0$. The metric supplementary equation (II), gives in the limit $\lambda\rightarrow 0$
$0=V_{(1),w}+\epsilon\left[-V_{(1)}^{2}+\kappa|\Phi_{(0),w}|^{2}\right]$ (76)
@@
This equation can be solved along the null boundary $\mathcal{B}^{\epsilon}$ for $w>0$, provided an initial value $V_{(1)}(0)$ is given at $w=0$ and provided the function $\Phi_{(0),w}$ is known for all values $w\geq 0$. The function $V_{(1)}$ measures the inaffinity of the null geodesics $n^{a}\partial_{a}$ ruling the null hypersurface $\mathcal{B}^{\epsilon}$, which follows from (12). For a non-vanishing and $w$-dependent scalar field on $\mathcal{B}^{\epsilon}$, (76) shows that the evolution of the inaffinity along $\mathcal{B}^{\epsilon}$ is driven by the free function $\Phi_{(0),w}$. Thus, the function $\Phi_{(0),w}$ can be seen as some kind of analogue of the news function at $\lambda=0$ instead of in the asymptotic limit $\lambda\rightarrow\infty$.
@@
The characteristic initial value problem can now be constructed in the following way. Let $w=0$ be a null hypersurface $\mathscr{N}^{\epsilon}(0)$ whose generators are tangent to affinly parameterized null geodesics with tangent $\ell^{a}\partial_{a}$ and affine parameter $\lambda$. At the origin of the affine parameter, $\lambda=0$, the null hypersurface $\mathscr{N}^{\epsilon}(0)$ intersects a null hypersurface $\mathcal{B}^{\epsilon}$ which has the generators $n^{a}\partial_{a}$ that obey the normalization $n^{a}\ell_{a}=-1$. The intersection of the two null hypersurfaces $\mathscr{N}^{\epsilon}(0)$ and $\mathcal{B}^{\epsilon}$ is a two dimensional spacelike space with spherical topology and unit Gaussian curvature, called $\Sigma$. The intrinsic metric of $\Sigma$ is the one of the unit sphere, $q_{AB}$. For this double null characteristic initial value problem, the following specification of data determines a solution of the Einstein-scalar field equations in spherical symmetry:
@@
  • specify the type of null hypersurface $w=0$, by choosing either $\epsilon=1$ (ingoing null hypersurface) or $\epsilon=-1$ (outgoing null hypersurface)
  • specify the gauge charge $q$
@@
  • on $\mathscr{N}^{\epsilon}(0)$ specify a free complex function $\mathcal{F}_{\Phi}(\lambda)$ that can be expressed by a regular Taylor series at $\lambda=0$
  • on $\mathcal{B}^{\epsilon}$ specify a free complex function $\mathcal{N}_{\mathcal{B}}(w)$
  • on $\Sigma$ specify the following fields: $\Phi(0,0)$ and $V_{(1)}(0)$
  • on $\mathcal{B}^{\epsilon}$, i.e. for $\lambda=0$, the following fields are trivial: $r$, $V$, $Q$, $\alpha$, $Z$ and $\mathcal{L}$
  • on $\mathcal{B}^{\epsilon}$, the field $r_{,\lambda}$ is nontrivial, i. e. $\lim_{\lambda\rightarrow 0}r_{,\lambda}=1$ for all values $w\geq 0$
With these specifications the field equations are solved in the following way:
@@
    with $\mathcal{N}_{\mathcal{B}}(w)$ and $V_{(1)}(0)$ determine the inaffinity everywhere on $\mathcal{B}^{\epsilon}$ via the equation (which follows from (76))
@@
    with $\mathcal{F}_{\Phi}(\lambda)$ and $\Phi(0,0)$ determine the scalar field everywhere on $\mathscr{N}^{\epsilon}$ via the equation
@@
    as all boundary data for $r$, $Q$, $\alpha$, $Z$, $\mathcal{L}$ and $V$ are known on $\mathcal{B}^{\epsilon}$, the hypersurface equations (29a) - (29f) can be hierarchically integrated for all the fields relevant to determine the $w$-derivative of $\Phi$ in equation (29g)
@@
    integration of the $w$-derivative of $\Phi$ by equation (29g) (using for example a finite difference scheme) determines $\Phi(\Delta w,\lambda)$ at some time $w=\Delta w>0$ on a null hypersurface $\mathscr{N}^{\epsilon}(\Delta w)$
@@
This algorithm can be suitably implemented in a numerical algorithm.

### v3

@@ III.2 Local initial-boundary formulation at a world line
In spherical symmetry there exists a designated world line, which is the geodesic tracing the centers of spherical symmetry Synge (1960). We denote this geodesic with $c(w)$ and choose $w$ as an affine parameter for its parameterization.
The tangent vector $c^{a}\partial_{a}$ is a unit timelike vector and $w$ measures the proper time along $c(w)$. This proper time is assumed to be the proper time of a freely falling Fermi observer along $c(w)$ Manasse and Misner (1963); Poisson (2009). Following Madler and Muller (2013), in a convex normal neighborhood of $c(w)$, we extend the definition of $w$ to be constant along null cones emanating from $c(w)$. The surface forming rays of the $w=$const null cones are labeled by the angular coordinates $x^{A}$ within each null cone. The null condition on $w=$const implies $g^{ww}=0$, since $g^{ab}\nabla_{a}w\nabla_{b}w=0$. The covector $\ell_{a}=\epsilon\nabla_{a}w$ defines by metric duality the tangent vectors $\ell^{a}=\epsilon g^{ab}\nabla_{b}w=\epsilon g^{aw}$ of the surface forming null rays. The null rays are labeled by the angular coordinates, $\ell^{a}\nabla_{a}x^{A}=0$ and therefore $g^{wA}=0$.
By these coordinate conditions, only the coordinate $x^{1}=\lambda$ varies along the rays of a given null cone $w=$const. Since the cones emanate from $c(w)$, we take the origin $\lambda=0$ to be on $c(w)$. We normalize the parameter $\lambda$ along the rays such that it is an affine parameter with the normalization determined at $c(w)$. In particular, $c^{a}\nabla_{a}\lambda=\ell^{a}\nabla_{a}\lambda=\epsilon g^{w\lambda}=1$ with $|\epsilon|=1$ at $c(w)$ so that $\epsilon=-1$ designates that the respective null cone $w=$const are outgoing null cones with null rays pointing into the future of $c(w)$ and $\epsilon=1$ are ingoing null cones pointing into the past of $c(w)$ (see Fig. 2).
Figure 2: A family of null cones $\mathscr{N}^{\epsilon}(w)$ at the central geodesic $c(w)$ is shown for $\epsilon=-1$ (left) and $\epsilon=+1$ (right). The left graphic shows the local geometry for the characteristic initial value problem where initial data are provided in the null cone $\mathscr{N}^{-}(0)$ and the right graphic illustrates it for the characteristic past value problem with the future data on $\mathscr{N}^{+}(0)$. The origin of the affine parameter $\lambda=0$ for each null cone $\mathscr{N}^{-}(w)$ is at the vertices of the cones traced by $c(w)$. Distinctive vertices are labeled with different values of $w$.
By these conditions, the metric near $c(w)$ has the form Madler and Muller (2013) [^7] Reference Madler and Muller (2013) uses a reversed nomenclature for the binary parameter $\epsilon$.
@@
The supplementary equation (II) holds at $c(w)$ due to the regularity conditions. Equations (63) and (18) imply
$Q_{,w}=O(\lambda^{2})\;\;\Rightarrow\;\;Q(w,0)=Q_{,\lambda}(w,0)=0.$ (65)
@@
where the subscript $(k)$ in parentheses indicate that the symbol carrying the subindex is the $\lambda^{k}$ coefficient function depending on $w$ of the Taylor series of the respective field evaluated at $\lambda=0$.
The scalar field expansions (66) and (29h) give
@@
Insertion of these data into (29) and integration give
@@
We can see that the time evolution of the coefficient $\Phi_{(i)}$ along $c(w)$ is driven by a function depending on the expansion coefficients containing all expansion coefficients up to $\Phi_{(i+1)}$. This behavior is a result of the choice of $w$ being constant along a null hypersurface. The hierarchical structure of (69) is the scalar field equivalent to the similar structure in the gravitational case Madler and Muller (2013); Mädler (2013) for gravitational waves. Since $\Phi_{(1)}=\lim_{\lambda\rightarrow 0}\Phi_{,\lambda}$ and $\Phi_{(2)}=\lim_{\lambda\rightarrow 0}\Phi_{,\lambda\lambda}$, we learn from (68) that the evolution of the initial data $\Phi$ along $c(w)$ and in its neighborhood is completely determined by the data on a cone $w=$const, provided the data and its derivatives are finite in the limit $\lambda\rightarrow 0$. Indeed, it is not difficult to derive similar equations like in (69) for $\Phi_{(k)}$ with $k\geq 2$, while using an appropriate higher order expansion than (66) and integrating (29).
From such an exercise, we can deduce the following local CIBVP at the vertices of null cones emanating from a geodesic world line: Suppose, the complex scalar field $\Phi(w,\lambda)$ on a null cone $w=0$ has a regular Taylor series expansion,
@@
evaluated at the world line, the boundary conditions of the fields are (63) and (64), $\alpha_{0}$ is an arbitrary function of $w$ for $w\neq 0$, and the gauge charge $q$ is provided; then, the metric, the electromagnetic potential $\alpha$ and the charge function $Q$ can be completely calculated on the cone $w=0$ using the hypersurface equations (29a)-(29f). Their local solution at $\lambda=0$ is of the form (68a), (68b), (68c), or (68f). The evolution equations (29g) determine the time derivative of the Taylor coefficients of the initial data, as seen in, e.g., (69).
We remark, that this regularity assumption at $w=0$ does not imply that the data necessarily remain regular for $w\neq 0$. If $r\rightarrow 0$ for $\lambda\rightarrow\lambda_{C}>0$ a (massive) caustic forms at finite value for the affine parameter. This caustic formation for some $w=w_{0}\neq 0$ is in general driven by the value of the norm squared of the derivative of the scalar field, $|\Phi_{,\lambda}|^{2}$. An example for caustic initial data was recently given in Mädler et al. (2024), where the data were employed in an evolution through an event horizon of a supercritical scalar field. The formulated initial value problem along a world line for a spherically symmetric charged scalar field is an analog to the gravitational CIBVP Friedrich (1986).
@@ III.3 Initial-boundary value problem at a null hypersurface
Consider a family of null hypersurfaces $w=$const designated by $\mathscr{N}^{\epsilon}(w)$, hereafter. The generators $\ell^{a}\partial_{a}$ of these null hypersurfaces are affinely parametrized with the affine parameter $\lambda$. We choose the origin, $\lambda=0$, of the affine parameter to be at a null hypersurface $\mathcal{B}^{\epsilon}$. The surface $\mathcal{B}^{\epsilon}$ is a null boundary for every null hypersurface $w=$const of $\mathscr{N}^{\epsilon}$ and is the boundary surface for the double-null CIBVPs to be presented, see Fig.(3).
Figure 3: A family of null hypersurfaces $\mathscr{N}^{\epsilon}(w)$ with corresponding boundary null hypersurface $\mathcal{B}^{\epsilon}(w)$ is shown for $\epsilon=-1$ (left) and $\epsilon=+1$ (right). The former is a characteristic initial value problem with the initial data surface $\mathscr{N}^{-}(0)$ and the latter gives rise to a characteristic past value problem with the past data on $\mathscr{N}^{+}(0)$. The common intersection between $\mathscr{N}^{\epsilon}(0)$ and $\mathcal{B}^{\epsilon}(w)$ is $\Sigma(0)$.
@@
This implies $V(w,0)=0$ for all values of $w$, where $\lambda=0$. Consequently, $V=O(\lambda)$ near $\mathcal{B}^{\epsilon}$ on $\mathscr{N}^{\epsilon}$. For the conformal factor $r$ there are two possibilities regarding its evaluation on $\mathcal{B}^{\epsilon}$: (i) $r(w,0)=0$, and (ii) $r(w,0)\neq 0$.
Physically these cases mean that the null hypersurface $\mathcal{B}^{\epsilon}$ has either a trivial [case (i)] or nontrivial Misner-Sharp mass [case (ii)], which becomes clear considering (21) in the limit $\lambda\rightarrow\infty$
@@ III.3.1 Initial boundary formulation at a null hypersurface with $r(w,0)=0$
If $V(w,0)=r(w,0)=0$, the line element near the origin of the affine parameter $\lambda$ can be written as
@@
Consequently, we can set wlog $r_{(1)}=1$, which is assumed hereafter in this subsection. With the assumption of a regular expansion of the scalar field as in (66), we find that $Z(w,0)=\mathcal{L}(w,0)=0$. The residual gauge freedom of the electromagnetic field $\alpha$ allows us to set $\alpha(w,0)=0$. In addition, the supplementary equation (18) implies $Q(w,0)$=const for $\lambda=0$ and we set wlog $Q(w,0)=0$. The supplementary equation (II) gives in the limit $\lambda\rightarrow 0$
$0=V_{(1),w}+\epsilon\left[-V_{(1)}^{2}+\kappa|\Phi_{(0),w}|^{2}\right].$ (76)
@@
This equation can be solved along the null boundary $\mathcal{B}^{\epsilon}$ for values $w\neq 0$, provided an initial value $V_{(1)}(0)$ is given at $w=0$ and provided the function $\Phi_{(0),w}$ is known for all values $w\geq 0$. The function $V_{(1)}$ measures the inaffinity of the null geodesics $n^{a}\partial_{a}$ generating the null hypersurface $\mathcal{B}^{\epsilon}$, which follows from (12). For a nonvanishing and $w$-dependent scalar field on $\mathcal{B}^{\epsilon}$, (76) shows that the evolution of the inaffinity along $\mathcal{B}^{\epsilon}$ is driven by the free function $\Phi_{(0),w}$.
Thus, the function $\Phi_{(0),w}$ can be seen as some kind of analog of the news function at $\lambda=0$ instead of in the asymptotic limit $\lambda\rightarrow\infty$.
@@
A CIBVP can now be constructed in the following way. Let $w=0$ be a null hypersurface $\mathscr{N}^{\epsilon}(0)$ whose generators are tangent to affinely parametrized null geodesics with tangent $\ell^{a}\partial_{a}$ and affine parameter $\lambda$. At the origin of the affine parameter, $\lambda=0$, the null hypersurface $\mathscr{N}^{\epsilon}(0)$ intersects a null hypersurface $\mathcal{B}^{\epsilon}$ which has the generators $n^{a}\partial_{a}$ that obey the normalization $n^{a}\ell_{a}=-1$. The intersection of the two null hypersurfaces $\mathscr{N}^{\epsilon}(0)$ and $\mathcal{B}^{\epsilon}$ is a two-dimensional spacelike space with spherical topology and unit Gaussian curvature, called $\Sigma$. The intrinsic metric of $\Sigma$ is the one of the unit sphere, $q_{AB}$. For this double-null CIBVP, the following specification of data determines a solution of the Einstein-scalar field equations in spherical symmetry:
@@
  • specify the type of null hypersurface $w=0$, by choosing either $\epsilon=1$ (ingoing null hypersurface) or $\epsilon=-1$ (outgoing null hypersurface);
  • specify the gauge charge $q$;
@@
  • on $\mathscr{N}^{\epsilon}(0)$ specify a free complex function $\mathcal{F}_{\Phi}(\lambda)$ that can be expressed by a regular Taylor series at $\lambda=0$;
  • on $\mathcal{B}^{\epsilon}$ specify a free complex function $\mathcal{N}_{\mathcal{B}}(w)$;
  • on $\Sigma$ specify the following fields: $\Phi(0,0)$ and $V_{(1)}(0)$;
  • on $\mathcal{B}^{\epsilon}$, i.e. for $\lambda=0$, the following fields are trivial: $r$, $V$, $Q$, $\alpha$, $Z$ and $\mathcal{L}$;
  • on $\mathcal{B}^{\epsilon}$, the field $r_{,\lambda}$ is nontrivial, i. e. $\lim_{\lambda\rightarrow 0}r_{,\lambda}=1$ for all values $w\geq 0$.
With these specifications, the field equations are solved in the following way:
@@
    With $\mathcal{N}_{\mathcal{B}}(w)$ and $V_{(1)}(0)$ determine the inaffinity everywhere on $\mathcal{B}^{\epsilon}$ via the equation [which follows from (76)]
@@
    With $\mathcal{F}_{\Phi}(\lambda)$ and $\Phi(0,0)$ determine the scalar field everywhere on $\mathscr{N}^{\epsilon}(0)$ via the equation
@@
    As all boundary data for $r$, $Q$, $\alpha$, $Z$, $\mathcal{L}$ and $V$ are known on $\mathcal{B}^{\epsilon}$, the hypersurface equations (29a) - (29f) can be hierarchically integrated for all the fields relevant to determine the $w$-derivative of $\Phi$ in Eq. (29g)
@@
    Integration of the $w$-derivative of $\Phi$ by Eq. (29g) (using, for example, a finite difference scheme) determines $\Phi(\Delta w,\lambda)$ at some time $w=\Delta w\neq 0$ on a null hypersurface $\mathscr{N}^{\epsilon}(\Delta w)$
@@
This algorithm can be suitably implemented numerically.